graph the equation and identify the y intercept 2y+5x=-2
Answer by rfer(16322)

2y+5x=-2
2(0)+5x=-2
x=-2/5
(-2/5,0) x int
------------
2y+5(0)=-2
2y=-2
y=-1
(0,-1) y int